A point charge pf -6 x 10^-9 C is 5 m from point A and 3 m from point B.

Find the Potential at point A and point B. What is the potential difference between

point A to B?


1
Expert's answer
2022-03-09T10:32:57-0500

The potential due point charge

"V=k\frac{q}{r}"

"V_1=9*10^9*\frac{6*10^{-9}}{5}=11\:\rm V"

"V_2=9*10^9*\frac{6*10^{-9}}{3}=18\:\rm V"

The potential difference between point A to B

"V=V_2-V_1=7\:\rm V"